# The Evolution and Internal Structure of Neptunes and Sub-Neptunes

### v1

@@ Abstract
Context. Neptunes and sub-Neptunes are typically modeled under the assumption that the interior is adiabatic and consists of distinct layers. However, this assumption is oversimplified as formation models indicate that composition gradients can exist. Such composition gradients can significantly affect the planetary thermal evolution by inhibiting convection. In non-convective layers, the heat transport is governed by multiple processes each relevant in different regions within the planet.
@@
Methods. We simulate the planetary evolution with an appropriate implementation for the conductivity by considering thermal transport via radiation, electrons, and vibrational conductivity. We consider planetary masses of 5, 10 and 15 M, three different initial energy budgets, and two different primordial composition profiles.
Results. We find that the assumed conductivity significantly affects the planetary thermal evolution. We show that the commonly-used conductivity assumption is inappropriate for modeling this planetary type. Furthermore, we find that the inferred radii deviate by $\sim 20\%$ depending on the assumed conductivity. The uncertainty on the primordial entropy in planets with non-convective layers leads to a difference of $\sim 25\%$ in the radii. This shows that the theoretical uncertainties are significantly larger than the observed ones, and emphasizes the importance of these parameters.
@@ 1 Introduction
Intermediate-mass planets make up the majority of the detected exoplanet population. Their interior is assumed to be composed mostly of rocks and ices surrounded by a hydrogen-helium (H-He) envelope. Unlike super-Earth planets, their radii are inflated by the presence of the light elements (e.g. Rogers et al., 2011; Lozovsky et al., 2018). The mass fraction of H-He strongly influences the radius regardless of the internal energy or the external stellar flux. Apart from the existence of a H-He atmosphere and a deep interior that is dominated by heavy elements, their detailed structure is highly uncertain. Constrained by only radius and mass, the composition profile is highly degenerate. Even for Uranus and Neptune, where gravitational data provide additional constraints on the planetary density distribution, the structure and bulk composition is still uncertain (e.g. Nettelmann et al., 2013; Neuenschwander & Helled, 2022; Morf et al., 2024).
@@
Since the planetary internal structure can change as the planet evolves, static models may be further constrained by modeling the planetary evolution. Often, planets are assumed to be fully convective, i.e., adiabatic, which imposes strong assumptions on both their evolution and structure. However, recently it became evident that assuming homogeneous and adiabatic structure is oversimplified. Composition gradients, which inhibit convection, can form due to the continuous accretion of solids and gas (e.g. Valletta & Helled, 2022) or by accreting different materials when migrating over ice-lines in the protoplanetary disc (e.g. Schneider & Bitsch, 2021; Eberlein et al., 2024). Indeed, it has been suggested that both Uranus and Neptune have composition gradients in their deep interiors that can survive for billions of years (Vazan & Helled, 2020; Tejada Arevalo, 2025). Composition gradients are not always primordial and can be caused by other processes such as phase separation and rainout of heavy elements. This further complicates the picture and can create (additional) composition gradients during the evolution (e.g., Stevenson, 1980; Lorenzen et al., 2009; Redmer et al., 2009; Püstow et al., 2016; Schöttler & Redmer, 2018; Chang et al., 2024; Cano Amoros et al., 2024). In addition, depending on the opacity, a deep radiative layer can also form within the planet (e.g., Howard et al., 2023; Müller & Helled, 2024).
@@
In non-convective regions, the heat transport is governed by multiple thermal transport mechanisms - radiative, diffusive, and conductive. Usually, the diffusive transport where hot particles diffuse into regions with colder temperatures is insignificant compared to radiation and conduction. In the outer parts of the planet, it is mostly the photons that contribute to thermal transport. Material emits and reabsorbs thermal radiation, and given a temperature gradient, this leads to a net diffusion of thermal energy that depends on the Rosseland-mean opacity (Freedman et al., 2014; Marigo & Aringer, 2009; Marigo et al., 2024). In denser regions, where the temperature is still sufficiently low so that the material is not ionized, conductivity is governed by the vibration of molecules, similar to phonons in solids. Ross et al. (1984) reviewed the phonon model for metals applied to anharmonic non-metallic crystals. This model was used for the thermal conductivity of the Earth’s Core-Mantle-Boundary (Stamenković et al., 2011). More recently, French (2019) used ab initio calculations of water to create an empirical fit of the thermal conductivity contribution from the nuclei. In the deep, hot interior of the planet, free electrons enhance the thermal conductivity. Cassisi et al. (2007) elaborated the chemical picture of the electron conductivity by considering the density and scattering processes of electrons. French & Redmer (2017) complemented the conductivities of the water nucleus by calculating the electron contribution.
@@
In this paper, we investigate the influence of the thermal conductivity on the evolution of Neptunes and sub-Neptunes. We compare different cases for the conductivity to assess how it affects the radius evolution and the inferred internal structure. Our paper is structured as follows. In Section 2, we summarize the physical processes and the methods. In section 3 we show the results of our study and discuss them in section 4. Finally, our summary and conclusions are presented in section 5.
@@ 2 Methods
We model the planetary evolution by solving the stellar structure equations (e.g. Kippenhahn et al., 2013) using the Henyey method (Henyey et al., 1965) with the Modules for Experiments in Stellar Astrophysics (MESA) code (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). We modified the code to make it applicable to medium-sized planets. We use the equation of state (EoS) presented by Müller et al. (2020b, a); Müller & Helled (2021, 2024) that combines the CMS EoS with non-ideal interaction (Chabrier & Debras, 2021) for H-He with the heavy-element EoS for H2O and SiO2 (More et al., 1988; Vazan et al., 2013) in a 50/50 rock-water mixture. The tables are implemented using a modified version of the MESA extension ’custom EoS’ (Knierim & Helled 2024; Helled et al. submitted). The modifications enable the solver to utilize linear interpolations within the EoS tables, rather than the standard bicubic spline interpolation scheme employed by MESA.
Furthermore, we change the radiative opacities and thermal conductivities as discussed below. We start the evolution with a given initial thermal state, composition profile and mass as discussed in section 2.2 and evolve the planet for 10 Gyr. The atmospheric cooling is simulated using the irradiated gray implementation in MESA based on Guillot & Havel (2011). We adopt an equilibrium temperature of $T_{\text{eq}}=400$ K and a mean visible-to-thermal opacity ratio of $\kappa_{\text{V}}/\kappa_{\text{th}}=0.03$, based on the fit provided by Poser & Redmer (2024). Although our chosen equilibrium temperature lies outside the original fit range of 500–4000 K, the authors estimate the fit uncertainty to be approximately $\pm 0.1$, which exceeds the difference in the opacity ratio between 400 K and 500 K. We use the Ledoux criterion to determine if a region is convective or non-convective. Because of numerically challenging simulations, we do not change the composition over time.
@@ 2.1 Thermal conductivity and opacity
A temperature gradient $\nabla T$ within a planet creates a thermal flux $F_{\text{thermal}}$. The flux within the planetary interior can be modeled as a diffusive process such that it follows Fourier’s law:
@@
given the thermal conductivity $k$. Then the total conductivity $k_{\text{tot}}$:
@@
is the sum of all processes where $\kappa_{\text{rad}}$ is the conductivity contribution from photons, and $k_{\text{vib}}$ and $k_{\text{elec}}$ are the conductivity contributions from vibrations in a dense fluid/solid and from electrons, respectively.
The radiative conductivity can be converted to radiative opacity $\kappa_{\text{rad}}$ which measures how much light is absorbed in a medium using:
@@
where $\sigma$ is the Stefan-Boltzmann constant, $T$ the temperature, and $\rho$ the density. In the following, we use the term radiative conductivity to refer to thermal transport by photons. The standard low temperature opacity tables in MESA that cover high metallicities are an extended version of the Freedman opacities (Freedman et al., 2008, 2014). They are tabulated for a fixed solar H-He ratio and different metallicity fractions up to $Z=1$, relevant for planets. However, the reliable range stops at $\log(T/\text{K})=3.8$, and significantly underestimates the opacity. In this region MESA uses high temperature opacities instead, which do not cover high $Z\geq 0.2$ materials. To improve on that, we created tables similar to the Freedman tables using the AESOPUS2.1 (Marigo & Aringer, 2009; Marigo et al., 2024) web interface. The recently updated method allows us to the create tables in the range for temperatures up to $\log(T/\text{K})=4.5$ with arbitrary values for the metallicity. The AESOPUS2.1 opacities were compared to the Freedman tables and to another set of opacities from Malygin et al. (2014). In the regime relevant for planets (high $\log R$), the AESOPUS2.1 opacities are generally higher than the opacities from the other two studies, especially below temperatures of $\log(T/\text{K})\lesssim 3.3$ for Malygin et al. (2014) and $\log(T/\text{K})\lesssim 3.5$ for the Freedman tables. The differences might originate from a more complete list of molecular lines over the entire temperature region (Marigo et al., 2024). As a result, the AESOPUS2.1 opacities seem to be more reliable. Note that the latest release of MESA includes a set of opacities tables made with the AESOPUS2.1 code. However, these tables only include low metallicities and are therefore inappropriate for modeling a large range of planetary types. For details on how we created the tables and used them in MESA, see the Appendix A.
@@
For the vibrational conductivity, we use two models and compare them with each other. From the theory of phonons in metals, Ross et al. (1984) reviewed the application to anharmonic non-metallic crystals. We use the formulation from Stamenković et al. (2011) and take the reference conductivity for MgSiO3 $k_{\text{vib, ref}}(\rho_{\text{ref}}=3.87\text{ g/cm}^{3},\,T_{\text{ref}}=2000\text{ K})=0.71$ W/m/K (see Table 1 Stamenković et al., 2011). More details can be found in Appendix B. The second model we use is based on the empirical fit to ab initio calculated conductivities of water (eq. 7 in French 2019). They provide an analytical function that is suitable for densities in the range of $\rho\in(0.2,10)$ g/cm3 and temperatures of $T\in(600,\,50\,000)$ K.
@@
The electron contribution to the thermal conductivity in MESA is modeled by default with improved tables from Cassis et al. 2007. This conductivity assumes fully ionized matter, which is not suitable for planets. We compare this conductivity with the fit to the ab initio calculations for the electron conductivity of partially ionized water (eq. 30 in French & Redmer, 2017). Their fit is valid in the same temperature density region as the work in French (2019).
@@
Because MESA uses the opacity formulation of the thermal transport, we combine the contributions from vibrations and electrons as a sum of the conductivities and convert the result into an opacity. This opacity is then combined with the radiative opacity using the reciprocal sum to get the total opacity. To avoid truncation errors in the reciprocal sum due to very low conductivities, we impose a lower bound of $k\geq 10^{-6}$ W/m/K on all the conductivity values. This threshold has little effect on the planetary evolution since it is at least five orders of magnitude smaller than the lowest conductivity obtained in any of our models. To verify this, we performed a simulation with a reduced lower limit of $k\geq 10^{-8}$ W/m/K and found a deviation of only $\sim 10^{-9}$ R in radius after 10 Gyr.
@@ 2.2 Model setups
Table 1: Used models for the conductivity by vibration and electrons. All models use the AESOPUS2.1 radiative opacity.
@@
We create models with a wide and narrow composition gradient. Except for Section 3.2, we use the wide transition profile if not stated otherwise. The envelope metallicity is fixed to $Z_{\text{env}}=0.20$ and the deep interior has $Z_{\text{int}}=1.0$. The transition region has a width of $\Delta q=0.1$ in terms of the normalized mass coordinate $q$ for the wide gradient and $\Delta q=0.01$ for the narrow gradient. The position is fixed to achieve a bulk composition of $Z_{\text{bulk}}=0.95$ with a hydrogen-helium ratio of $Y/X=0.34$ based on the solar photosphere (Asplund et al., 2009). Figure 1 shows the initial specific entropy, composition, and temperature profile for the wide composition gradient (see Figure 11 for the narrow composition gradient). Details on how we use MESA’s relax options to create our initial model are presented in Appendix C.
@@
To compare the different approaches to the conductivity, we create multiple model cases: Cond-1 uses the vibrational and electronic conductivity for water under high pressure (French & Redmer, 2017; French, 2019), Cond-2 uses the default electron conductivity of MESA for fully ionized material based on Cassisi et al. (2007), Cond-3 assumes a constant electron+vibration conductivity based on the conductivity at the Earth’s Core-Mantle-Boundary of $k_{\text{elec}}=4$ W/m/K (e.g. Stevenson et al., 1983; Lobanov et al., 2021), and Cond-4 uses the phonon model for the vibrational conductivity (Stamenković et al., 2011) and the electron conductivity for partially ionized water (French & Redmer, 2017). All the cases use the AESOUPS tables for the radiative conductivity. We summarize the models in Table 1. Overall, we recommend using Cond-1 because it is the most updated model. However, currently the conductivity of Cond-1 is only available for H2O and we hope that future work investigates the thermal conductivity of various materials under a large range of pressures and temperatures.

### v2

# arXiv:2509.04564v2[astro-ph.EP] 02 Mar 2026

@@ Abstract
Context. Neptunes and sub-Neptunes are typically modeled under the assumption that the interior is adiabatic and consists of distinct layers. However, this assumption is oversimplified, as formation models indicate that composition gradients can exist. Such composition gradients can significantly affect the planetary thermal evolution by inhibiting convection. In non-convective layers, the heat transport is governed by multiple processes, and each is relevant in different regions within the planet.
@@
Methods. We simulated the planetary evolution of such objects with an appropriate implementation for the conductivity by considering thermal transport via radiation, electrons, and vibrational conductivity. We considered planetary masses of 5, 10, and 15 M; three different initial energy budgets; and two different primordial composition profiles.
Results. We find that the assumed conductivity significantly affects the planetary thermal evolution. We show that the commonly used conductivity assumption is inappropriate for modeling this planetary type. Furthermore, we find that the inferred radii deviate by $\sim 20\%$ depending on the assumed conductivity. The uncertainty on the primordial entropy in planets with non-convective layers leads to a difference of $\sim 25\%$ in the radii. This shows that the theoretical uncertainties are significantly larger than the observed ones and emphasizes the importance of these parameters.
@@ 1 Introduction
Intermediate-mass planets make up the majority of the detected exoplanet population. Their interiors are assumed to be composed of mostly rocks and ices surrounded by a hydrogen-helium (H-He) envelope. Unlike super-Earth planets, their radii are inflated by the presence of light elements (e.g. Rogers et al., 2011; Lozovsky et al., 2018). The mass fraction of H-He strongly influences the radius of these planets regardless of the internal energy or the external stellar flux. Apart from the existence of an H-He atmosphere and a deep interior that is dominated by heavy elements, their detailed structure is highly uncertain. Constrained by only radius and mass, the composition profile of the detected exoplanet population is highly degenerate. Even for Uranus and Neptune, where gravitational data provide additional constraints on the planetary density distribution, the structure and bulk composition is still uncertain (e.g. Nettelmann et al., 2013; Neuenschwander & Helled, 2022; Morf et al., 2024).
@@
Since the internal planetary structure can change as a planet evolves, static models may be further constrained by modeling planetary evolution. Often, planets are assumed to be fully convective, i.e., adiabatic, which imposes strong assumptions on both their evolution and structure. However, it recently became evident that assuming a homogeneous and adiabatic structure is an oversimplification. Composition gradients, which inhibit convection, can form due to the continuous accretion of solids and gas (e.g. Valletta & Helled, 2022) or by accreting different materials when migrating over ice lines in the protoplanetary disk (e.g. Schneider & Bitsch, 2021; Eberlein et al., 2024). Indeed, it has been suggested that both Uranus and Neptune have composition gradients in their deep interiors that can survive for billions of years (Vazan & Helled, 2020; Tejada Arevalo, 2025). Composition gradients are not always primordial and can be caused by other processes, such as phase separation and the rainout of heavy elements. This further complicates the picture and can create (additional) composition gradients during the evolution of a planet (e.g., Stevenson, 1980; Lorenzen et al., 2009; Redmer et al., 2009; Püstow et al., 2016; Schöttler & Redmer, 2018; Chang et al., 2024; Cano Amoros et al., 2024). In addition, depending on the opacity, a deep radiative layer can also form within the planet (e.g., Howard et al., 2023; Müller & Helled, 2024).
@@
In non-convective regions, the heat transport is governed by multiple thermal transport mechanisms – radiative, diffusive, and conductive. Usually, the diffusive transport where hot particles diffuse into regions with colder temperatures is insignificant compared to radiation and conduction. In the outer parts of the planet, it is mostly the photons that contribute to thermal transport. Material emits and reabsorbs thermal radiation, and given a temperature gradient, this leads to a net diffusion of thermal energy that depends on the Rosseland-mean opacity (Freedman et al., 2014; Marigo & Aringer, 2009; Marigo et al., 2024). In denser regions, where the temperature is still sufficiently low so that the material is not ionized, conductivity is governed by the vibration of molecules, similar to phonons in solids. Ross et al. (1984) reviewed the phonon model for metals applied to anharmonic non-metallic crystals. This model was used for the thermal conductivity of the Earth’s core-mantle boundary (Stamenković et al., 2011). More recently, French (2019) used ab initio calculations of water to create an empirical fit of the thermal conductivity contribution from the nuclei. In the deep hot interior of the planet, free electrons enhance the thermal conductivity. Cassisi et al. (2007) elaborated on the chemical picture of the electron conductivity by considering the density and scattering processes of electrons. French & Redmer (2017) complemented the conductivities of the water nucleus by calculating the electron contribution.
@@
In this paper, we investigate the influence of thermal conductivity on the evolution of Neptunes and sub-Neptunes. We compare different cases for the conductivity to assess how it affects the radius evolution and the inferred internal structure. Our paper is structured as follows. In Section 2, we summarize the physical processes and the methods. In section 3 we show the results of our study and discuss them in section 4. Finally, our summary and conclusions are presented in section 5.
@@ 2 Methods
We modeled the planetary evolution by solving the stellar structure equations (e.g. Kippenhahn et al., 2013) using the Henyey method (Henyey et al., 1965) with the Modules for Experiments in Stellar Astrophysics (MESA) code (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). We modified the code to make it applicable to medium-sized planets. We used the equation of state (EoS) presented by Müller et al. (2020b, a); Müller & Helled (2021, 2024) that combines the EoS with non-ideal interaction (Chabrier & Debras, 2021) for H-He with the heavy-element EoS for H2O and SiO2 (More et al., 1988; Vazan et al., 2013) in a 50/50 rock-water mixture. The tables were implemented using a modified version of the MESA extension ’custom EoS’ (Knierim & Helled, 2024; Helled et al., 2025). The modifications enabled the solver to utilize linear interpolations within the EoS tables rather than the standard bicubic spline interpolation scheme employed by MESA.
Furthermore, we changed the radiative opacities and thermal conductivities as discussed below. We started the evolution with a given initial thermal state, composition profile, and mass as discussed in section 2.2 and evolved the planet for 10 Gyr. The atmospheric cooling was simulated using the irradiated gray implementation in MESA based on Guillot & Havel (2011). We adopted an equilibrium temperature of $T_{\text{eq}}=400$ K and a mean visible-to-thermal opacity ratio of $\kappa_{\text{V}}/\kappa_{\text{th}}=0.03$ based on the fit provided by Poser & Redmer (2024). Although our chosen equilibrium temperature lies outside the original fit range of 500–4000 K, the authors estimate the fit uncertainty to be approximately $\pm 0.1$, which exceeds the difference in the opacity ratio between 400 K and 500 K. We used the Ledoux criterion to determine if a region is convective or non-convective. Because of numerically challenging simulations, we did not change the composition over time.
@@ 2.1 Thermal conductivity and opacity
A temperature gradient, $\nabla T$, within a planet creates a thermal flux, $F_{\text{thermal}}$. The flux within the planetary interior can be modeled as a diffusive process such that it follows Fourier’s law:
@@
given the thermal conductivity, $k$. The total conductivity, $k_{\text{tot}}$, which is calculated as
@@
is the sum of all processes, where $\kappa_{\text{rad}}$ is the conductivity contribution from photons, while $k_{\text{vib}}$ and $k_{\text{elec}}$ are the conductivity contributions from vibrations in a dense fluid and from electrons, respectively.
The radiative conductivity can be converted to radiative opacity, $\kappa_{\text{rad}}$, which measures how much light is absorbed in a medium using
@@
where $\sigma$ is the Stefan-Boltzmann constant, $T$ the temperature, and $\rho$ the density. In the following, we use the term radiative conductivity to refer to thermal transport by photons. The standard low temperature opacity tables in MESA that cover high metallicities are an extended version of the Freedman opacities (Freedman et al., 2008, 2014). They have been tabulated for a fixed solar H-He ratio and different metallicity fractions up to $Z=1$, relevant for planets. However, the reliable range stops at $\log(T/\text{K})=3.8$ and significantly underestimates the opacity. In this region, MESA uses high temperature opacities instead, which do not cover high $Z\geq 0.2$ materials. To improve on that, we created tables similar to the Freedman tables using the AESOPUS2.1 (Marigo & Aringer, 2009; Marigo et al., 2024) web interface. The recently updated method allowed us to the create tables in the range for temperatures up to $\log(T/\text{K})=4.5$ with arbitrary values for the metallicity. The AESOPUS2.1 opacities were compared to the Freedman tables and to another set of opacities from Malygin et al. (2014). In the regime relevant for planets (high $\log R$), the AESOPUS2.1 opacities are generally higher than the opacities from the other two studies, especially below temperatures of $\log(T/\text{K})\lesssim 3.3$ for Malygin et al. (2014) and $\log(T/\text{K})\lesssim 3.5$ for the Freedman tables. The differences might originate from a more complete list of molecular lines over the entire temperature region (Marigo et al., 2024). As a result, the AESOPUS2.1 opacities seem to be more reliable. We note that the latest release of MESA includes a set of opacity tables made with the AESOPUS2.1 code. However, these tables only include low metallicities and are therefore inappropriate for modeling a large range of planetary types. (For details on how we created the tables and used them in MESA, see the Appendix A.)
@@
For the vibrational conductivity, we used two models and compared them with each other. From the theory of phonons in metals, Ross et al. (1984) reviewed the application to anharmonic non-metallic crystals. We used the formulation from Stamenković et al. (2011) and took the reference conductivity for MgSiO3 $k_{\text{vib, ref}}(\rho_{\text{ref}}=3.87\text{ g/cm}^{3},\,T_{\text{ref}}=2000\text{ K})=0.71$ W/m/K (see Table 1 Stamenković et al., 2011). More details can be found in Appendix B. The second model we used is based on the empirical fit to ab initio calculated conductivities of water (eq. 7 in French 2019). They provide an analytical function that is suitable for densities in the range of $\rho\in(0.2,10)$ g/cm3 and temperatures of $T\in(600,\,50\,000)$ K.
@@
The electron contribution to the thermal conductivity in MESA is modeled by default with improved tables from Cassis et al. (2007). This conductivity assumes fully ionized matter, which is not suitable for planets. We compared this conductivity with the fit to the ab initio calculations for the electron conductivity of partially ionized water (eq. 30 in French & Redmer, 2017). Their fit is valid in the same temperature density region as the work in French (2019).
@@
Because MESA uses the opacity formulation of the thermal transport, we combined the contributions from vibrations and electrons as a sum of the conductivities and converted the result into an opacity. This opacity was then combined with the radiative opacity using the reciprocal sum to get the total opacity. To avoid truncation errors in the reciprocal sum due to very low conductivities, we imposed a lower bound of $k\geq 10^{-6}$ W/m/K on all the conductivity values. This threshold has little effect on the planetary evolution since it is at least five orders of magnitude smaller than the lowest conductivity obtained in any of our models. To verify this, we performed a simulation with a reduced lower limit of $k\geq 10^{-8}$ W/m/K and found a deviation of only $\sim 10^{-9}$ R in radius after 10 Gyr.
@@ 2.2 Model setups
Table 1: Models used for the conductivity by vibration and electrons. All models use the AESOPUS2.1 radiative opacity.
@@
We created models with a wide and narrow composition gradient. Except for what we report in Section 3.2 and if not otherwise stated, we used the wide transition profile. The envelope metallicity was fixed to $Z_{\text{env}}=0.20$ and the deep interior has $Z_{\text{int}}=1.0$. The transition region has a width of $\Delta q=0.1$ in terms of the normalized mass coordinate $q$ for the wide gradient and $\Delta q=0.001$ [^1] The published version contained a typo, stating the width of the narrow gradient as $\Delta q=0.01$. The correct value used in the simulations is $\Delta q=0.001$. for the narrow gradient. The position was fixed to achieve a bulk composition of $Z_{\text{bulk}}=0.95$ with a hydrogen-helium ratio of $Y/X=0.34$ based on the solar photosphere (Asplund et al., 2009). Figure 1 shows the initial specific entropy, composition, and temperature profile for the wide composition gradient (see Figure 11 for the narrow composition gradient). Details on how we used MESA’s relax options to create our initial model are presented in Appendix C.
@@
To compare the different approaches to the conductivity, we created multiple model cases: Cond-1 uses the vibrational and electronic conductivity for water under high pressure (French & Redmer, 2017; French, 2019), Cond-2 uses the default electron conductivity of MESA for fully ionized material based on Cassisi et al. (2007), Cond-3 assumes a constant electron plus vibration conductivity based on the conductivity at the Earth’s core-mantle boundary of $k_{\text{elec}}=4$ W/m/K (e.g. Stevenson et al., 1983; Lobanov et al., 2021), and Cond-4 uses the phonon model for the vibrational conductivity (Stamenković et al., 2011) and the electron conductivity for partially ionized water (French & Redmer, 2017). All the cases use the AESOUPS tables for the radiative conductivity. We summarize the models in Table 1. Overall, we recommend using Cond-1 because it is the most updated model. However, the conductivity of Cond-1 is currently only available for H2O, and we hope that future work investigates the thermal conductivity of various materials under a large range of pressures and temperatures.